OpenStudy (anonymous):

10 = x divided by 4 minus 8 10=x/4-8

OpenStudy (saifoo.khan):

\[10 = \frac x4 -8\]this?

OpenStudy (saifoo.khan):

\[\frac x4 = 18\]cross multiply. \[x = 18 * 4\]x = 72

OpenStudy (anonymous):

i got x= -40 because i did 10=(x/4-8) first i did the subtraction so now it looks like this 10=(x/-4) then i multiplied the -4 on both sides (to get x alone) and i got x=-40

OpenStudy (anonymous):

10=( x/4)-8 ==> x/4=18 ==> x=4*18 =72

OpenStudy (anonymous):

\[10=(x/4) -8\] \[x/4 = 18\] x= 72